OpenStudy (anonymous):

condense the expression: 1/2Log5 16-3log5 x+4Log5y

OpenStudy (anonymous):

first move the coefficients that are infront of the logarithms to the right

OpenStudy (anonymous):

meaning log\[\log_{5} 16^{1/2}\]

OpenStudy (anonymous):

ignore the first log my equations button is not working

OpenStudy (anonymous):

but the log expression below is correct use that example to simplify the other two

OpenStudy (anonymous):

then let me know what you got

OpenStudy (anonymous):

log5y^x+4?

OpenStudy (anonymous):

\[\log_{5} 4-\log_{5} x ^{3}+\log_{5} y ^{4}\]

OpenStudy (anonymous):

thankyou for helping me their just really hard for me ;(

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

have time for more?

OpenStudy (anonymous):

we aren't done yet sorry, now when you are subtract two logs with the same base do you condense them by multiplying them together or dividing them?

OpenStudy (anonymous):

dividing

OpenStudy (anonymous):

yes so log base 4 is on top the numerator making the other two...?

OpenStudy (anonymous):

on the bottom being multiplied together

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

because they were being added

OpenStudy (anonymous):

so can you show me what your answer would be?

OpenStudy (anonymous):

4/x^3+y^4

OpenStudy (anonymous):

when you are adding two logs together with the same base then you multiply them together

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[\log_{5} (4/x ^{3}y ^{4})\]

OpenStudy (anonymous):

does that make sense

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i just want to make sure you understand rather than just giving you the answer so that later you can ace a test on this if you need to

OpenStudy (anonymous):

yah this good

OpenStudy (anonymous):

ok why we try another problem. this will be the last one i can help with, then i have to go alright?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

solve the equation; 15+2 log2 x=31

OpenStudy (anonymous):

so when solving for any variable the whole point is to get the variable by itself so what would be one thing we can move away from the side with x easily?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

that was a question

OpenStudy (anonymous):

and the answer is 15

OpenStudy (anonymous):

woudnt it be 17?

OpenStudy (anonymous):

i didn't mean the final answer, i meant that it will be easy to move the fifteen since the fifteen is not attached to anything

OpenStudy (anonymous):

move to be 15

OpenStudy (anonymous):

yes we subtract fifteen from both sides then we simplify the logarithm but first what does the new equation look like?

OpenStudy (anonymous):

it will equeal 16instead of 31

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

know we focus on the log

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

okay!

OpenStudy (anonymous):

what can we move to condense the logarithm more?

OpenStudy (anonymous):

the 2 in front

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ah yes!!

OpenStudy (anonymous):

so what do we have now?

OpenStudy (anonymous):

logx=16

OpenStudy (anonymous):

no... \[\log_{2} x ^{2}=16\]

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay a trick that i learned is you put the same base of the log you want to remove under both sides of the equation like so: \[2^{\log_{2}x ^{2} }=2^{16}\]

OpenStudy (anonymous):

now what can you simplify?

OpenStudy (anonymous):

exponents

OpenStudy (anonymous):

yes \[x ^{2}=2^{16}\] is our new equation as 2^log base 2 equals one

OpenStudy (anonymous):

now what would be the next step?

OpenStudy (anonymous):

logx=4

OpenStudy (anonymous):

why would we do that when we already have x by itself. what we would do is simplify 2^16=65536

OpenStudy (anonymous):

then x^2=65536

OpenStudy (anonymous):

65536?

OpenStudy (anonymous):

yes now what would you do to get x all by itself?

OpenStudy (anonymous):

take the root of x and the other side

OpenStudy (anonymous):

x=256

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

thats what i thought

OpenStudy (anonymous):

\[\sqrt{x ^{2}}=\sqrt{65536}=x=256\]

OpenStudy (anonymous):

omg thankyou so much

OpenStudy (anonymous):

no problem I have to go now, but i wish you the best of luck :)

OpenStudy (anonymous):

thankyou